OpenStudy (anonymous):

Indicate which property is illustrated in Step 8. Step 1 4 - 7x - 4 + 8x = (4 - 7x) + (-4 + 8x) Step 2 = (4 - 7x) + (8x - 4) Step 3 = 4 + (-7x + 8x) - 4 Step 4 = 4 + (-7 + 8)x - 4 Step 5 = 4 + x - 4 Step 6 = x + 4 - 4 Step 7 = x + 0 Step 8 = x identity distributive commutative associative

OpenStudy (anonymous):

@Directrix its not helping at all

Directrix (directrix):

x + 0 TO x Adding zero doesn't change anything. Which property? @Starr_Dynasty

Directrix (directrix):

@Starr_Dynasty >>>> its not helping at all Don't despair.

OpenStudy (anonymous):

Indicate which property is illustrated in Step 4. Step 1 10 + 8x - 10 - 2x + 2 = 10 + (8x - 10) - 2x + 2 Step 2 = 10 + (-10 + 8x) - 2x + 2 Step 3 = [10 + (-10)] + (8x - 2x) + 2 Step 4 = 0 + (8x - 2x) + 2 Step 5 = (8x - 2x) + 2 Step 6 = (8 - 2)x + 2 Step 7 = 6x + 2 associative additive inverse commutative distributive

Directrix (directrix):

[10 + (-10)] + (8x - 2x) + 2 TO 0 + (8x - 2x) + 2

OpenStudy (anonymous):

could i just have the answer

Directrix (directrix):

If I knew it, maybe. But, I'm having to figure this out.

Directrix (directrix):

[10 + (-10)] + (8x - 2x) + 2 TO 0 + (8x - 2x) + 2

Directrix (directrix):

This is what happened" [10 + (-10)] went to 0

Directrix (directrix):

That is some kind of inverse. Do you have an inverse option? @Starr_Dynasty

Directrix (directrix):

@Starr_Dynasty Did you get this?

Directrix (directrix):

END OF THREAD. BEGIN A NEW ONE. THANKS.